# From Code to Play: Benchmarking Program Search for Games Using Large Language Models

### v1

@@ V-F Tabletop Games Framework (TAG)
  • Related to this, they are also inherently multiplayer. As such there is implicit opponent modeling required for good play strategies. The environment is no longer a ‘simple’ stationary MDP, but is actively adversarial.
  • The TAG framework has a number of local libraries and coding conventions; for example decks of cards are implemented via Deck<> or PartialObservableDeck<> parameterised classes. These are not likely to be present in the LLM training data to any degree, and require the LLM to generalise to unseen software architecture details. This contrasts to the straightforward Python with mostly standard libraries of the games in earlier sections.
  • The language used is now Java. Integration of all language models used langchain4j. [^12] https://docs.langchain4j.dev
Algorithm 1 was applied to 12 tabletop board games (see Table VIII for the full list) implemented in TAG. These are multi-player environments (2 to 4 players) with partial observability and stochasticity and varying levels of complexity.
TABLE VIII: TAG results by game. Key as for Table IX, plus P is the number of players, Best Agent records the model that won the round robin tournament. SM is the number of models that produced working code on at least one iteration and entered an agent in the round robin tournament. BB indicates if the best agent significantly Beats the Baseline agent (OSLA or MCTS); $\approx$ means performance matches the baseline.
Game P S.Iter. SM Best Agent BB
Can’t Stop 3 66% 8 GPT 4o Yes
Colt Express 3 28% 7 Claude Haiku $\approx$
Connect 4 2 34% 7 Claude Haiku $\approx$
Diamant 4 30% 5 Claude Sonnet No
Dominion 3 35% 6 GPT 4o Yes
Hearts 4 24% 5 Claude Sonnet Yes
Love Letter 3 29% 8 GPT 4o $\approx$
Poker 4 35% 7 Gemini Pro Yes
Seven Wonders 4 16% 5 Claude Sonnet Yes
Sushi Go! 4 16% 6 Gemini Pro No
Tic-Tac-Toe 2 26% 6 Gemini Pro Yes
Virus 2 59% 7 Mistral Small No
Given the additional level of complexity of these games, and very different (and often dynamic) action spaces for each game, the language models were not tasked with writing a full policy to play the game. Instead they were tasked with writing a heuristic function to estimate the value of a game state for a player. This should be close to 1.0 for a position that is a definite win, to 0.0 for a position that is a definite loss. This heuristic function was then used within a search algorithm; either one step lookahead (OSLA) or Monte Carlo Tree Search (MCTS).
@@
    Automatic rulebook digestion. This takes as input the PDF of the game rulebook. An approach inspired by [52] is used. The rulebook is first broken down into chunks of 1000 or 2000 words. The LLM is then given each chunk in turn and asked to summarise in 200 words or less the information about the game rules. This final set of synopses is then fed to the LLM with a prompt to, ‘Summarise this information in 500 words or less.’. This provides an blocks of text to include in the prompt used in the main loop of Algorithm 1 that explains the rules of the game.
@@
The multi-player nature of these environments also necessitates a change in the evaluation criterion in Algorithm 1. Evaluation used a tournament of 500 games. The base agent in the tournament was either a one step lookahead (OSLA) agent (for Tic-Tac-Toe and Virus) or a vanilla MCTS agent (all other games) with a budget of 10ms and a rollout of 10 actions before the generated heuristic function is applied.
@@
A base opponent used a heuristic function of the game score (for all other games). All games in Table VIII have the concept of score except for Tic-Tac-Toe and Connect 4, which only reward a win (+1) or draw (+0.5). To avoid overfitting to a specific opponent, later iterations within a run of Algorithm 1 add all previous (working) agents to the evaluation tournament. The evaluation score of each generated heuristic is the win rate from the most recent tournament, so this is updated to include a broader range of opponents later in the run.
For each game a final tournament is then run between the best agents from each model, for a maximum of 8 participants if all models generated at least one heuristic that compiled and executed successfully. Model performance is then judged by giving 5 points to the winner of each tournament, 4 points to the second-place and down to 1 point for 5th place.
@@
TABLE IX: TAG results by model. S.Iter is the percentage of iterations that resulted in working code. Cost is the total cost of all 120 iterations (10 per game) on the LLM and Points are from round robin tournaments between the best agents from each LLM for each game. For each game 5 points are given for first place, 4 for 2nd and so on down to 1 for 5th place. Zero points are awarded otherwise, including for LLMs that failed to produce any working code for a game. The maximum number of points is therefore 60.
Model S.Iter. Cost (\$) Points
Claude Sonnet 38% 11.81 35
Claude Haiku 14% 1.18 15
Gemini Pro 48% 3.22 31
Gemini Flash 25% 0.24 14
Mistral Large 33% 3.06 20
Mistral Small 19% 0.51 18
GPT 4o 48% 6.95 31
GPT 4o mini 40% 0.30 16
Table IX summarises the results by language model. Large models do consistently better than their smaller counterpart in terms of both the number of successful iterations (ones that generate java code that compiles and runs) and in the quality of the best heuristics produced. However the smaller models are much cheaper to run, and do sometimes generate winning code. This suggests that given a fixed budget for LLM calls using a smaller model for a larger number of iterations and runs could be the better choice, although this has not been tested here experimentally.
There is no major difference between the four model families. GPT and Gemini fare slightly better overall, especially in terms of generating more consistently valid code; but this is in aggregate and there is high variability across games. The larger GPT model failed to produce a single instance of working code for Virus for example (while the large Claude and Gemini models both had 9 or 10 functional iterations for the same task and prompt). Similarly, on three games the large Gemini model only had a single iteration out of the 10 that generated working code.
The small Anthropic model (Claude 3.5 Haiku) is poor overall, but still produces the single best performing heuristic in 2 of the 12 games. This variability is actually very encouraging as gives traction to this semi-evolutionary approach. If each model gave the same response to the same prompt then we could not make progress (all models use their default temperature and TopP or TopK settings). There is no clear pattern in terms of which model is better at which type of game, and these results suggest that using a variety of models to generate even wider phenotypic variety can be beneficial.
@@
Overall results by game are shown in Table VIII. One common reason for failure of an iteration was code that compiled but then failed to execute in all edge cases due to poor error checking for division by zero (throwing a runtime error during the evaluation tournament was counted as a failure of the iteration). Otherwise the models were often quite creative in their invention of undocumented API methods causing compilation to fail.
The complexity of API to an LLM is not always the same as complexity of a game. Tic-Tac-Toe and Connect 4 are simple games that are won by building a straight line of a player’s markers (3 and 4 respectively). Each is represented by a grid of cells. In TAG this is represented by a parameterised GridBoard<T extends Token> class, with methods that require navigating a deeper class hierarchy and understanding parameterised types. There is a int getPlayerAt(int x, int y) helper method on the game state as well, but the LLMs often fail to write code that compiles due to issues with GridBoards. The game with the highest iteration success rate, Can’t Stop, is also the only game for which the base game state has no dependencies outside the core java.lang and java.util libraries.
@@
Sushi Go! had a particularly low success rate for generating valid code. A specific problem here was that the card types in the game were represented by an enum that had values expressed in lower case; Maki, Sashimi, Dumpling. Despite these values being clearly stated in the Java API section of the prompt generated code more commonly used MAKI, SASHIMI, DUMPLING. This is presumably because an upper case convention is more standard across Java more generally and hence in the training data of the models.
These results only compare the agents from the language model results with each other. In a few cases the ‘best’ agent was not actually very good at the game, and was beaten by an OSLA/MCTS agent using the game score as very simple heuristic as in [53]. This is shown in Table VIII. Six of the games (i.e. half) having a heuristic generated (across 80 iterations in total) that could reliably beat the baseline score heuristic, and three fail to generate one that is even as good as simply targeting the score.
VI Conclusion
@@
In this work we studied and evaluated the current possibilities of using LLMs for program search in the area of games for various applications. Previous work was mostly limited to a single problem or game without being easily transferable to other domains, as the DSL had to be adapted. We demonstrated that LLMs can overcome the problem of combinatorial explosion of search spaces constructed with predefined DSLs, and that LLMs are able to synthesize programmatic policies in Python for the Minatar domain, which was not possible with a custom DSL and previous methods. Furthermore, we have shown that this framework can be easily adapted to different applications by modifying the prompts, and that it often provides reasonable results even without much customization. We have shown that even with the default temperature settings on these standard language models there is a very wide range of output for the same input prompt; in this respect at least the models can be quite ‘creative’. Running many independent iterations of the same task can create a varied population of outputs. This is very promising as it provides the variation required for the evolutionary hill-climbing approach used here.
@@
We observed limitations in the quality of the generated code. For example, in the simple 2D vehicle driving task, the generated code drove the car to the target but then failed to stop most of the time. Much of the generated code fails to run, or in the case of Java, to compile. These limitations become more evident as the complexity of the task increases. The need to use framework-specific Java libraries in TAG leads to less than 1 in 5 attempts generating valid code. We believe limitations such as this could be overcome with more sophisticated search and better prompt engineering, but the results so far give an idea of the limitations of what can be achieved with relatively little effort. The addition of tools to the LLM interfaces and a more agentic workflow is a promising area for this future work. For example instead of asking the LLM to generate the code in one pass, it could be asked to construct useful component functions or sub-modules with documented interfaces. In a later pass the model could then be asked to combine these sub-modules (based on feedback of performance of previous combinations).
One general recommendations that can be made is to use a variety of models as there can be huge variability between them on a given task. Given a robust evaluation method, as in the 29 tasks here, the smaller models can be much more cost effective at up to a tenth of the cost of their larger brethren and just as able to produce top quality results if given the time.
VII Acknowledgements

### v2

@@ V-G Tabletop Games Framework (TAG)
  • They are also inherently multiplayer. As such there is implicit opponent modeling required for good play strategies. The environment is no longer a ‘simple’ stationary MDP, but is adversarial.
  • The TAG framework has a number of local libraries and coding conventions; for example decks of cards are implemented via Deck<> or PartialObservableDeck<> parameterised classes. These are not likely to be present in the LLM training data, and require the LLM to generalise to unseen software architecture details. This contrasts to standard Python with common libraries of the games in earlier sections.
  • The language used is now Java. Integration of all language models apart from Llama used langchain4j [^12] https://docs.langchain4j.dev. Llama experiments used the Gemini Vertex AI interface to access the Google Llama model-as-a-service [^13] https://console.cloud.google.com/vertex-ai/publishers/meta/model-garden.
Algorithm 2 was applied to 12 tabletop board games (see Table VIII) implemented in TAG. These are adversarial multi-player environments with partial observability and stochasticity and varying levels of complexity. One player count in the supported range for each game was selected for use to give an even distribution of player counts between 2 and 4.
TABLE VIII: TAG results by game. S.Iter is the percentage of iterations that resulted in working code and S.Trl the percentage of Trials (each consisting of 10 iterations). P is the number of players, Best Agent records the model that won the round robin tournament. SM is the number of models that produced any working code and entered an agent in the round robin tournament. BB indicates if the best agent significantly Beats the Baseline agent (OSLA or MCTS); $\approx$ means performance matches the baseline.
Game P S.Iter S.Trl SM Best Agent BB
Can’t Stop 3 65% 91% 11 o1-mini Yes
Colt Express 3 19% 44% 10 Llama 3.3 70B Yes
Connect 4 2 61% 88% 11 o1-mini $\approx$
Diamant 4 39% 89% 11 GPT 4o mini Yes
Dominion 3 13% 44% 7 Gemini Flash Yes
Hearts 4 31% 65% 10 GPT 4o Yes
Love Letter 3 19% 37% 9 o1-mini $\approx$
Poker 4 18% 50% 11 GPT 4o mini Yes
Seven Wonders 4 25% 40% 7 GPT 4o Yes
Sushi Go! 4 30% 58% 10 Gemini Flash Yes
Tic-Tac-Toe 2 67% 87% 11 No differences Yes
Virus 2 29% 55% 8 Gemini Flash Yes
Given the additional level of complexity of these games, and different (and often dynamic) action spaces for each game, the language models were not asked to write a full policy to play the game. Instead they were asked to write a heuristic function to estimate the value of a game state for a player. This should be close to 1.0 for a position that is a definite win, to 0.0 for a position that is a definite loss (other heuristics are possible, taking account of relative scores or ordinal positions, but this simple win/loss estimate keeps things simple). This heuristic function was then used within a search algorithm; either one step lookahead (OSLA) or Monte Carlo Tree Search (MCTS) [53].
@@
    Automatic rulebook digestion. This takes as input the text in the game rulebook. An approach inspired by [54] is used. The PDF rulebook is first broken down into chunks of 1000 or 2000 words (so any icons or pictures are excluded). The LLM is then given each chunk in turn and asked to summarise in 200 words or less the information about the game rules. This final set of synopses is then fed to the LLM with a prompt to, ‘Summarise this information in 500 words or less.’. This provides blocks of text to include in the prompt used in the main loop of Algorithm 2 that explains the rules of the game.
@@
The multi-player nature of these environments also necessitates a change in the evaluation criterion in Algorithm 2. Evaluation used a tournament of 500 games between the new agent and a base agent. The base agent was either a one step lookahead (OSLA) agent (for Tic-Tac-Toe and Virus) or a vanilla MCTS agent (all other games) with a budget of 10ms and a rollout of 10 actions before the generated heuristic function is applied to estimate the value of the state. TAG uses Multiplayer-UCT with the heuristic applied for each player independently [55]. This small budget enables the large number of experiments to be run in a reasonable time, but will not give the best possible players, although we go get perfect play in Tac-Tac-Toe at about 30ms with no heuristic. It remains an open question how this might change with a larger budget, but what is important here is the comparative performance of the generated heuristics.
@@
A base opponent used the same OSLA or MCTS settings and a heuristic function of the game score normalised to [0, 1]. Tic-Tac-Toe and Connect 4 do not have scores, and the base opponents for these rewarded a win (+1) or draw (+0.5), with 0 for a loss. To avoid overfitting to a specific opponent all previous (working) agents to the evaluation tournament are added to later iterations within a trial. The evaluation score of each generated heuristic is the win rate from the most recent tournament, so this includes a broader range of opponents later in the trial. This performance metric means that, unlike the single-player environments, the scores from each run are not directly comparable. This modifies lines 12-17 of Algorithm 2; EvaluateFitness returns the current best-performing of the agents, and this is retained as the new bestResult. Each trial is restarted with just the base opponent.
For each game a final tournament of 25,000 games is run between the best agents from each model, for a maximum of 11 participants if all models generated at least one heuristic that compiled and executed successfully. Points in Table IX are awarded for 1st through 10th places in this final tournament for each game. This methodology compares the results of the models directly and rewards an LLM that produces a variety of heuristics, including some that are very good at the game, over one that reliably generates valid, but near-identical code each time.
@@
TABLE IX: TAG results by model. S.Iter and S.Trl are the same as for Table IX, plus FG is the number of games for which the LLM failed to produce any working code. Cost is the total cost of all 1200 iterations (100 per game) on the LLM. For each game in a round-robin tournament between the agents, 10 Points are given for the first place, 9 for 2nd, and so on down to 1 for 10th place. Zero points are awarded otherwise, including for LLMs that failed to produce any working code for a game. The Points column is the total of these, and Rank is the ranked order of models by points.
Model S.Iter S.Trl FG Cost (\$) Points Rank
Claude Sonnet 17% 82% 0 134.47 61 6
Claude Haiku 8% 50% 0 86.71 36 9
Gemini Flash 54% 78% 0 1.60 84 1st
Gemini Lite 28% 51% 0 2.93 63 5
Mistral Large 16% 34% 2 19.73 26 10
Mistral Small 17% 30% 7 3.24 20 11
o1-mini 58% 75% 1 23.92 69 3rd
GPT 4o 41% 62% 1 26.01 72 2nd
GPT 4o mini 32% 65% 1 4.99 67 4
Llama 3.3 70B 63% 82% 2 3.51 48 7
Llama 3.1 70B 49% 65% 1 0.00 41 8
Table IX summarises the results by language model. The Gemini 2.0 Flash model does best overall in the final tournament, generates valid code consistently and is much cheaper than the OpenAI models that also do very well. The older Mistral models do least well on these tasks on all measures; the larger Llama model is best at generating working code, but the performance of its best agents is relatively poor; the Anthropic models are much more expensive than the others, but this is not reflected in performance levels. The reason the smaller Gemini model is more expensive (it was 25% cheaper per token) is less working code was generated (S.Iter) so there were many more calls back to the model to fix compilation errors.
Large models do better on average than their smaller counterpart in terms of both the number of successful iterations, trials and in the quality of the best heuristics produced. However, in most cases the performance differences are small and this effect is smaller than the differences between the model families. GPT 4o mini is a fifth of the cost of GPT 40, but still writes the best agent in 2 games, and is only a few points behind in tournament points.
The fact that many trials fail to produce any working code over 10 iterations show the importance of re-starts, as using a single trial for each game led to more random results.
@@
Overall results by game are shown in Table VIII. One common reason for failure of an iteration was code that compiled but then failed to execute in all edge cases due to poor error checking for division by zero (throwing a runtime error during the evaluation tournament was counted as a failure of the iteration). There was no clear pattern of performance improvement across the 10 iterations of each trial. In some cases the later heuristics were better than the first attempts, but equally often the overall winner was the first heuristic found and later changes did not improve performance. The benefit of running more iterations for improved performance is investigated further in Section VI. The complexity of API to an LLM is not always the same as complexity of a game.
Otherwise the models were often creative in their invention of undocumented API methods causing compilation to fail. The game with the highest trial success rate, Can’t Stop, is also the only game for which the base game state has no dependencies outside the core java.lang and java.util libraries, reducing the opportunities for LLMs to hallucinate about non-existent methods.
Figure 6: The reward curves for each Minatar environment for the long-running experiments. The x-axis shows the number of iterations and the y-axis the reward of the best program. The dashed red-line shows the max. reward achieved by the best LLM from Table II. We use the same parameters as in Section V-A, only with 1000 iterations instead of 10 per trial.
@@
In contrast the card games in the set cause much more LLM code to fail to compile. All the 8 games in Table VIII with S.Iter below 35% use decks of cards, and none of the 4 games with S.Iter above this do. The TAG-specific infrastructure of Deck<> and PartialObservableDeck<> parameterised classes, with their own interfaces often caused problems. For example Deck.getSize() is used to return the number of cards, and despite this being mentioned in the prompt through recursion of the class hierarchy, many LLMs tried to call the non-existent Deck.size() or Deck.length(). A similar problem in Sushi Go! was that the card types in the game were represented by an enum that had values expressed in lower case; Maki, Sashimi, Dumpling. Despite these values being clearly stated in the Java API section of the prompt generated code more commonly used MAKI, SASHIMI, DUMPLING. This is presumably because an upper case convention is more standard across Java more generally and hence in the training data of the models.
Table VIII, column BB, shows that at least one of the LLMs wrote a heuristic that could reliably beat the baseline score heuristic in 10 of the 12 games, and was at least competitive in the other two. Running just one trial per game per LLM (10 iterations) does not reliably get baseline performance, and restarting is needed to get a good end result from multiple independent attempts.
VI A Thousand Iterations for Minatar
To evaluate the impact of the number of iterations on program search, we perform 1000 iterations for each game available in Minatar. We keep the same values for the other parameters, i.e. 10 trials for the program search and three attempts for generation/improvement or repair of non-executable programs. Figure 6 shows the reward curves for each environment for the long-running experiments. In most games (with the exception of Freeway and Asterix) there is a high variance in the rewards achieved in each trial, suggesting that more restarts are better than running for more iterations. The reward curves for Freeway support this statement, since in some trials, the programs found after early iterations are almost as good as after the last iterations. It is also clear that the max. reward of the best trial achieves a better result for Seaquest and Space Invader and approaches the same reward for Breakout as the best LLM from Table II. For Freeway, the LLM is not able to improve the reward compared to the 10 iterations experiment, while for Asterix the reward is almost doubled but still not as good as the best LLM from the previous experiments.
VII Practical Recommendations & Discussion
TABLE X: The ranking of the LLMs on the different domains with the overall ranking of all experiments. It is clear there are major differences between the domains.
LLM Minatar Maze Baba Vehicle TAG Overall
Claude Sonnet 3rd 6 1st 4 6 3rd
Claude Haiku 7 4 7 5 9 6
Gemini Flash 4 7 7 7 1st 4
Gemini Lite 11 5 7 9 5 10
Mistral Large 6 10 3rd 1st 10 5
Mistral Small 9 10 3rd 2nd 11 9
o1 mini 1st 1st 2nd 6 3rd 1st
GPT 4o 2nd 2nd 3rd 8 2nd 2nd
GPT 4o mini 5 2nd 11 10 4 6
Llama 3.3 70B 8 8 7 11 7 11
Llama 3.1 70B 10 9 3rd 3rd 8 8
Table X shows a global ranking of the LLMs for the different domains. It is clear that there are major differences between the model families; for example, Mistral models are best for vehicle driving, while the OpenAI models are best for generating mazes and the Minatar domain. o1 mini is overall the best model, which is expected since it is the only reasoning model, and thus uses more compute than the others. For generating code in Java for TAG, Gemini 2.0 Flash is the best model followed by the GPT model family. A rather unexpected result is that Llama 3.3 is worse than Llama 3.1, although Meta advertises that it achieves a performance comparable to the Llama 3.1 405B model. In general, the open-source Llama models cannot keep up with the closed-source models of the LLM providers. However, the reasons for this are difficult to analyse as we do not know the parameter size of the paid models and there are also differences in the training pipeline and dataset. The 70B-Llama models are probably larger than the small closed-source models, but much smaller than the large models. With enough iterations and restarts, the 70B model of Llama 3.1 achieves the same performance as the large models of OpenAI and Claude, but only with 100 times more iterations. This indicates that it is in principle possible to achieve the same performance with open-source models, they just need more time to find good programs.
After discussing the game applications and the long-running experiments, there are several practical recommendations we can make for using LLMs to synthesize programs for game research. These depend primarily on the available financial resources, hardware and time.
If money is no issue, it is best to try different models and to start with the larger models as good programs are found more quickly. With financial constraints, but local hardware to run the Llama 3.1 or 3.3 70B model, e.g. on a 48GB GPU with 4-bit quantization [56], then it is better to run local models for more iterations, as a similar reward to the Claude or OpenAI LLMs can be achieved. However, depending on the simulation environment, this may take much longer.
If there are financial and time constraints, we recommend using Table X to select the best models for a given domain.
One unexpected finding is that larger models are not always better, e.g. in the maze generation experiments, where most smaller models were better than their larger counterpart. Therefore, smaller models are worth trying and should not be dismissed from the start. They can also be much more cost-efficient overall even if they are run for many more trials/iterations. The results in Table IX make clear that total model costs can vary by 2 orders of magnitude with the cheaper model giving better results on a particular problem.
From the long-running experiments it is clear that more restarts are better than always running more iterations. The number of iterations depends heavily on the problem domain, and more difficult problems also need more iterations. This is visible in Figure 6 for Seaquest, the most complicated Minatar domain, where the reward increases only in later iterations compared to Freeway or Asterix.
VIII Conclusion
@@
In this work we studied and evaluated the current possibilities of using LLMs for program search in the area of games for various applications. Previous work was mostly limited to a single problem or game without being easily transferable to other domains, as the DSL had to be adapted. We demonstrated that LLMs can overcome the problem of combinatorial explosion of search spaces constructed with predefined DSLs, and that LLMs are able to synthesize programmatic policies in Python for the Minatar domain, which was not possible with a custom DSL and previous methods. Furthermore, we have shown that this framework can be easily adapted to different applications by modifying the prompts, and that it often provides reasonable results even without much customization. We have shown that even with the default temperature settings on these standard language models there is a very wide range of output for the same input prompt; in this respect at least the models can be quite ‘creative’. Running many independent iterations of the same task can create a varied population of outputs. This is very promising as it provides the variation required for the hill-climbing approach used here.
@@
We observed limitations in the quality of the generated code. For example, in the simple 2D vehicle driving task, the generated code drove the car to the target but then failed to stop most of the time. Much of the generated code fails to achieve any reward at all, or in the case of Java, to compile. These limitations become more evident as the complexity of the task increases. The need to use framework-specific Java libraries in TAG leads to less than 1 in 5 attempts generating valid code. We believe limitations such as this could be overcome with more sophisticated search and better prompt engineering, but the results so far give an idea of the limitations of what can be achieved with relatively little effort. The addition of tools to the LLM interfaces and a more agentic workflow is a promising area for this future work. For example instead of asking the LLM to generate the code in one pass, it could be asked to construct useful component functions or sub-modules with documented interfaces. In a later pass the model could then be asked to combine these sub-modules (based on feedback of performance of previous combinations).
IX Acknowledgements